# Scalable Bayesian Learning with posteriors

### v1

@@ Abstract
Although theoretically compelling, Bayesian learning with modern machine learning models is computationally challenging since it requires approximating a high dimensional posterior distribution. In this work, we (i) introduce posteriors111posteriors library: https://github.com/normal-computing/posteriors, an easily extensible PyTorch library hosting general-purpose implementations making Bayesian learning accessible and scalable to large data and parameter regimes; (ii) present a tempered framing of stochastic gradient Markov chain Monte Carlo, as implemented in posteriors, that transitions seamlessly into optimization and unveils a minor modification to deep ensembles to ensure they are asymptotically unbiased for the Bayesian posterior, and (iii) demonstrate and compare the utility of Bayesian approximations through experiments including an investigation into the cold posterior effect and applications with large language models.
@@ 1 Introduction
which provides a principled way to update our beliefs about the parameters $\theta\in\operatorname{\mathbf{R}}^{d}$ as data is observed once $y_{1:N}$ or sequentially $y_{N}$. Bayesian learning represents uncertainty about the parameters using the posterior distribution $p(\theta\mid y_{1:N})$, which combines prior beliefs $p(\theta)$ with the likelihood or generating process $p(y\mid\theta)$ for the data. This contrasts with optimization approaches, which store a single point estimate of the parameters $\theta^{*}=\operatorname*{arg\,max}_{\theta}p(y_{1:N}\mid\theta)$. In our view, for modern machine learning models and pipelines, Bayesian learning offers three key advantages, which are visualized in Figure 1: (a) improved generalization and out-of-distribution predictions, (b) coherent online learning of new information (without catastrophic forgetting), and (c) the ability to decompose predictive uncertainty into aleatoric (natural or explained uncertainty in the data) and epistemic (unexplained uncertainty that diminishes with more training data) components.
@@
Generalization is a key metric for machine learning models, quantifying the ability of the model to perform tasks beyond its training set. Bayesian predictions offer a compelling approach to mitigate overfitting by averaging over a posterior distribution of plausible models rather than relying on a single fit. Bayesian methods have demonstrated improved out-of-distribution performance [57, 8, 46].
@@
Another vital consideration for modern machine learning pipelines is that of online or continual learning, where the model continuously updates and learns from new data in real-time without needing to be retrained from scratch. Bayes’ theorem (1) provides a coherent framework for online learning, where new data can be sequentially incorporated into the model by using the posterior distribution as the prior on receipt of the next data batch. When implemented exactly, this is equivalent to learning all data $y_{1:N}$ at once (1), and all data points are exchangeable and treated identically. This contrasts to naive online optimization, which typically results in the catastrophic forgetting [27, 40] of previously learnt tasks.
@@
In addition, it is essential for statistical models to be auditable and reliable, especially when they are met with data that is far outside the training distribution. With an optimization approach, the model is only able to provide predictions in the form of total uncertainty for the predictive distribution over test inputs $x^{*}$, via the entropy $H[p(y\mid x,\theta^{*})]:=-\operatorname{\mathbb{E}}_{p(y\mid x,\theta^{*})}[\log p(y\mid x,\theta^{*})]$. This is in contrast to Bayesian methods, which can decompose the total uncertainty (TU) of the posterior predictive distribution $p(y\mid x)=\operatorname{\mathbb{E}}_{p(\theta\mid y_{1:N})}[p(y\mid x,\theta)]$ into two components [79, 71, 30]:
@@
Aleatoric uncertainty (AU) represents natural uncertainty in the data, such as synonyms or sentence starts in natural language generation. Epistemic uncertainty (EU) represents uncertainty in the model parameters and diminishes with more data (as $p(\theta\mid y_{1:N})\to\delta(\theta\mid\theta^{*})$ as $N\to\infty$) and therefore is a better indication of model uncertainty in predictions. This decomposition is crucial for understanding model behaviour and making informed decisions in high-stakes applications [38], although is not unique to Bayesian models but rather the broader concept of second-order uncertainty [61, 71, 30].
@@
Exact Bayesian learning is computationally challenging due to the intractability of the high-dimensional posterior distribution. This has led to the development of various approximate Bayesian learning methods, each with its own trade-offs in terms of flexibility, accuracy and computational cost. These methods have been implemented in many software packages e.g. [11, 6, 10]. Yet, these packages mostly focus on the traditional Bayesian setting of relatively small data and small parameter regimes and are not designed for the scale of many machine learning applications.
@@
    We introduce posteriors; a functional, minibatch-first PyTorch [62] package for scalable and extensible Bayesian learning.
@@
    We provide experiments demonstrating the broad utility of posteriors and the accessibility of the aforementioned benefits. We investigate the generalization performance of a host of approximate Bayesian methods and the prominence of the cold posterior effect [77, 35], composing posteriors with LoRA [31] to mitigate forgetting in a continual learning task with Llama 2 [74] and out-of-distribution detection via epistemic uncertainty with Llama 3 [1].
@@ 2 Bayesian learning
A Laplace approximation is motivated by a second-order Taylor expansion around the mode of the posterior distribution $\theta_{\text{MAP}}=\operatorname*{arg\,max}_{\theta}p(\theta\mid y_{1:N})$. This gives a Gaussian approximation $p(\theta\mid y_{1:N})\approx\operatorname{\mathbf{N}}(\theta\mid\theta_{\text{MAP}},\hat{\Sigma})$. In the supervised learning setting, the covariance $\hat{\Sigma}$ is typically chosen [15] to be an approximation to the Fisher information matrix $\operatorname{\mathrm{F}}(\theta)=\operatorname{\mathbb{E}}_{p(x,y\mid\theta)}[\nabla^{2}_{\theta}\log p(\theta\mid x,y)]$ which is intractable as we do not have access to the true distribution over inputs $p(x)$. Instead, two approximations [42] are used based on the training data $\hat{p}(x,y)=N^{-1}\sum_{i=1}^{N}\delta(x,y\mid x^{i},y^{i})$. The first we denote the empirical Fisher $\operatorname{\mathrm{F}}_{\text{E}}(\theta)=\operatorname{\mathbb{E}}_{\hat{p}(x,y)}[\nabla_{\theta}\log p(\theta\mid y)\nabla_{\theta}\log p(\theta\mid y)^{\top}]$ and is easy to compute. The second we denote the conditional Fisher $\operatorname{\mathrm{F}}_{\text{C}}(\theta)=\operatorname{\mathbb{E}}_{\hat{p}(x)p(y\mid x,\theta)}[\nabla_{\theta}\log p(\theta\mid y)\nabla_{\theta}\log p(\theta\mid y)^{\top}]$ and can be computed efficiently for common neural network likelihoods via an equivalence with the Generalized Gauss-Newton matrix (GGN) [54] with further details in A.1.
@@
In its most common form variational inference (VI) [7] also forms a Gaussian approximation $p(\theta\mid y_{1:N})\approx q_{\phi}(\theta)=\operatorname{\mathbf{N}}(\theta\mid\mu,\Sigma)$. However, VI directly learns the mean and covariance $\phi=(\mu,\Sigma)$ through an optimization routine to minimize the Kullback-Leibler objective $\text{KL}[q_{\phi}(\theta)\mid\mid p(\theta\mid y_{1:N})]$, whose gradient can be approximated with Monte Carlo samples and tricks including reparameterzation [67] and stick-the-landing [70].
@@
An important caveat to note for Gaussian approximations is that the full covariance matrix requires $\operatorname{\mathcal{O}}(d^{2})$ memory and often $\operatorname{\mathcal{O}}(d^{3})$ runtime for the required inversion and sampling operations. In the large parameter regime, this is prohibitive, and further approximations are required [50, 68, 15]. Additionally, we can reduce memory at inference time by constructing uncertainty directly in function space by linearizing the forward model [23, 33]. If the likelihood has the form $p(y\mid x,\theta)=p(y\mid f_{\theta}(x))$ and we have a posterior approximation $\operatorname{\mathbf{N}}(\theta\mid\mu,\Sigma)$ then the linearized forward distribution becomes $\operatorname{\mathbf{N}}(f(x)\mid f_{\mu}(x),\nabla_{\theta}f_{\mu}(x)^{\top}\Sigma\nabla_{\theta}f_{\mu}(x))$ where $\nabla_{\theta}f_{\mu}(x)$ is the Jacobian of $f_{\theta}(x)$ evaluated at $\mu$. This avoids having to sample many large parameter vectors.
@@
Stochastic gradient Markov chain Monte Carlo (SGMCMC) [48, 58] instead forms a Monte Carlo approximation $p(\theta\mid y_{1:N})\approx N^{-1}\sum_{k=1}^{K}\delta(\theta\mid\theta_{k})$. The samples $\{\theta_{k}\}_{k=1}^{K}$ are typically collected by (approximately) evolving a stochastic differential equation (SDE) - see Section 4 and Appendix B - and collecting samples along the trajectory. Traditional full-batch Markov chain Monte Carlo (MCMC) schemes [4] use a Metropolis-Hastings step to ensure convergence of the Monte Carlo approximation is preserved through discretization. In the minibatch setting, the Metropolis-Hastings ratio cannot be estimated easily. Instead, suppose we have access to an unbiased stochastic gradient. Then, as noted in [76], the bias in the trajectory decreases quadratically $\operatorname{\mathcal{O}}(\epsilon_{t}^{2})$ for learning rate $\epsilon_{t}$. Therefore, adopting a decreasing Robbins-Monro [69] learning rate schedule $\sum\epsilon_{t}^{2}<\sum\epsilon_{t}=\infty$ will ensure an asymptotically unbiased Monte Carlo approximation.
@@
  • Composability: posteriors is written in PyTorch [62] and has compatibility with a broad range of tools including the Llama [74] and Mistral [36] models, lightning [22] for convenient logging and device management, optimizers from torchopt [66] and probabilistic programs from pyro [6].
  • Extensible: The transform framework (Figure 3) adopted by posteriors is very general and allows for the easy adoption of new algorithms. Additionally, posteriors supports arbitrary likelihoods rather than being restricted to hard-coded regression or classification as is common in other libraries [15, 17].
  • Functional: posteriors adopts a functional API via PyTorch’s [62] functional module. As championed by the JAX [9] ecosystem, the functional approach makes for code that is easier to test and compose with other functions. Importantly for posteriors, functional programming is also closer to the mathematical description, which is particularly useful for Bayesian modelling.
  • Scalable: Unlike most common Bayesian software [11, 10, 6], posteriors is minibatch-first, thus allowing for efficient computation in large datasets. Additionally, flexible sub-spacing is provided for scaling to large models.
@@
In addition to the extensible suite of Bayesian methods, posteriors also supplies a host of general-purpose utility functions not found elsewhere in the PyTorch ecosystem. This includes linearized forward calls, Fisher/GGN matrices, their diagonal counterparts, fast Fisher/GGN vector products and the conjugate gradient method, which provides the tools to implement Hessian-free optimization [56] and experiment with other second-order methods [73, 20].
@@ 3.1 Limitations
posteriors supports arbitrary likelihoods and loss functions; this means that methods that utilise the specific form of the loss function (normally as classification or regression), e.g. the deterministic last layer approach in [28] or the probit approximations from [15], are currently out of scope. Similarly, posteriors does not assume any specific structure of the model, which makes model-aware techniques such as DVI [80] or KFAC [55] difficult to provide directly; however, the functional approach should allow easy composition with bespoke model-aware code. Finally, we opt not to include a fit convenience function [13, 15] in favour of the more flexible and transparent update framework in Figure 3.

### v2

@@ Abstract
Although theoretically compelling, Bayesian learning with modern machine learning models is computationally challenging since it requires approximating a high dimensional posterior distribution. In this work, we (i) introduce posteriors, an easily extensible PyTorch library hosting general-purpose implementations making Bayesian learning accessible and scalable to large data and parameter regimes; (ii) present a tempered framing of stochastic gradient Markov chain Monte Carlo, as implemented in posteriors, that transitions seamlessly into optimization and unveils a minor modification to deep ensembles to ensure they are asymptotically unbiased for the Bayesian posterior, and (iii) demonstrate and compare the utility of Bayesian approximations through experiments including an investigation into the cold posterior effect and applications with large language models.
posteriors: github.com/normal-computing/posteriors
@@ 1 Introduction
which provides a principled way to update our beliefs about the parameters $\theta\in\operatorname{\mathbf{R}}^{d}$ as data is observed once $y_{1:N}$ or sequentially $y_{N}$. Bayesian learning represents uncertainty about the parameters using the posterior distribution $p(\theta\mid y_{1:N})$, which combines prior beliefs $p(\theta)$ with the likelihood or generating process $p(y\mid\theta)$ for the data. This contrasts with optimization approaches, which store a single point estimate of the parameters $\theta^{*}=\operatorname*{arg\,max}_{\theta}p(y_{1:N}\mid\theta)$. In our view, for modern machine learning models and pipelines, Bayesian learning offers three key advantages, which are visualized in Figure 1: (a) improved generalization and out-of-distribution predictions, (b) coherent online learning of new information (without catastrophic forgetting), and (c) the ability to decompose predictive uncertainty into aleatoric (natural or explained uncertainty in the data) and epistemic (unexplained uncertainty that diminishes with more training data) components. It’s important to note that in the more classical setting of statistical models with interpretable parameters, there are additional benefits to be gained from Bayesian learning - these include propagation of uncertainty (Duffield et al., 2024) and model validation (Gelman et al., 2020). In this work, we introduce posteriors, a python package designed to be minibatch-first and easily applicable to large-scale models with PyTorch (Paszke et al., 2019) and Hugging Face (Wolf et al., 2020) whilst also being compatible with classical Bayesian models through composition with Pyro (Bingham et al. 2019, Appendix G).
@@
Generalization is a key metric for machine learning models, quantifying the ability of the model to perform tasks beyond its training set. Bayesian predictions offer a compelling approach to mitigate overfitting by averaging over a posterior distribution of plausible models rather than relying on a single fit. In various settings, Bayesian methods have demonstrated improved out-of-distribution performance (Neal, 2012; Blundell et al., 2015; Liu et al., 2020).
@@
Another vital consideration for modern machine learning pipelines is that of online or continual learning, where the model continuously updates and learns from new data in real-time without needing to be retrained from scratch. Bayes’ theorem (equation 1) provides a coherent framework for online learning, where new data can be sequentially incorporated into the model by using the posterior distribution as the prior on receipt of the next data batch. When implemented exactly, this is equivalent to learning all data $y_{1:N}$ at once equation 1, and all data points are exchangeable and treated identically. This contrasts to naive online optimization, which typically results in the catastrophic forgetting (Goodfellow et al., 2013; Kirkpatrick et al., 2017) of previously learnt tasks.
@@
In addition, it is essential for statistical models to be auditable and reliable, especially when they are met with data that is far outside the training distribution. With an optimization approach, the model is only able to provide predictions in the form of total uncertainty for the predictive distribution over test inputs $x^{*}$, via the entropy $H[p(y\mid x,\theta^{*})]:=-\mathbb{E}_{p(y\mid x,\theta^{*})}[\log p(y\mid x,\theta^{*})]$. This is in contrast to Bayesian methods, which can decompose the total uncertainty (TU) of the posterior predictive distribution $p(y\mid x)=\mathbb{E}_{p(\theta\mid y_{1:N})}[p(y\mid x,\theta)]$ into two components (Wimmer et al., 2023; Sale et al., 2023; Hofman et al., 2024):
@@
Aleatoric uncertainty (AU) represents natural uncertainty in the data, such as synonyms or sentence starts in natural language generation. Epistemic uncertainty (EU) represents uncertainty in the model parameters and diminishes with more data (as $p(\theta\mid y_{1:N})\to\delta(\theta\mid\theta^{*})$ as $N\to\infty$) and therefore is a better indication of model uncertainty in predictions. This decomposition is crucial for understanding model behaviour and making informed decisions in high-stakes applications (Kendall and Gal, 2017), although is not unique to Bayesian models but rather the broader concept of second-order uncertainty (Osband et al., 2024; Sale et al., 2023; Hofman et al., 2024).
@@
    We introduce posteriors; a functional, minibatch-first PyTorch (Paszke et al., 2019) package for scalable and extensible Bayesian learning.
@@
    We provide experiments demonstrating the broad utility of posteriors and the accessibility of the aforementioned benefits. We investigate the generalization performance of a host of approximate Bayesian methods and the prominence of the cold posterior effect (Wenzel et al., 2020; Izmailov et al., 2021), composing posteriors with LoRA (Hu et al., 2021) to mitigate forgetting in a continual learning task with Llama 2 (Touvron et al., 2023) and out-of-distribution detection via epistemic uncertainty with Llama 3 (AI@Meta, 2024).
@@ 2 Bayesian learning
A Laplace approximation is motivated by a second-order Taylor expansion around the mode of the posterior distribution $\theta_{\text{MAP}}=\operatorname*{arg\,max}_{\theta}p(\theta\mid y_{1:N})$. This gives a Gaussian approximation $p(\theta\mid y_{1:N})\approx\operatorname{\mathbf{N}}(\theta\mid\theta_{\text{MAP}},\hat{\Sigma})$. In the supervised learning setting, the covariance $\hat{\Sigma}$ is typically chosen (Daxberger et al., 2021) to be an approximation to the Fisher information matrix $\operatorname{\mathrm{F}}(\theta)=\mathbb{E}_{p(x,y\mid\theta)}[\nabla^{2}_{\theta}\log p(\theta\mid x,y)]$ which is intractable as we do not have access to the true distribution over inputs $p(x)$. Instead, two approximations (Kunstner et al., 2019) are used based on the training data $\hat{p}(x,y)=N^{-1}\sum_{i=1}^{N}\delta(x,y\mid x^{i},y^{i})$. The first we denote the empirical Fisher $\operatorname{\mathrm{F}}_{\text{E}}(\theta)=\mathbb{E}_{\hat{p}(x,y)}[\nabla_{\theta}\log p(\theta\mid y)\nabla_{\theta}\log p(\theta\mid y)^{\top}]$ and is easy to compute. The second we denote the conditional Fisher $\operatorname{\mathrm{F}}_{\text{C}}(\theta)=\mathbb{E}_{\hat{p}(x)p(y\mid x,\theta)}[\nabla_{\theta}\log p(\theta\mid y)\nabla_{\theta}\log p(\theta\mid y)^{\top}]$ and can be computed efficiently for common neural network likelihoods via an equivalence with the Generalized Gauss-Newton matrix (GGN, Martens 2020) with further details in Appendix A.1.
@@
In its most common form variational inference (VI, Blei et al. 2017) also forms a Gaussian approximation $p(\theta\mid y_{1:N})\approx q_{\phi}(\theta)=\operatorname{\mathbf{N}}(\theta\mid\mu,\Sigma)$. However, VI directly learns the mean and covariance $\phi=(\mu,\Sigma)$ through an optimization routine to minimize the Kullback-Leibler objective $\text{KL}[q_{\phi}(\theta)\mid\mid p(\theta\mid y_{1:N})]$, whose gradient can be approximated with Monte Carlo samples and tricks including reparameterzation (Rezende et al., 2014) and stick-the-landing (Roeder et al., 2017).
@@
An important caveat to note for Gaussian approximations is that the full covariance matrix requires $\operatorname{\mathcal{O}}(d^{2})$ memory and often $\operatorname{\mathcal{O}}(d^{3})$ runtime for the required inversion and sampling operations. In the large parameter regime, this is prohibitive, and further approximations are required (MacKay, 1992; Ritter et al., 2018; Daxberger et al., 2021). Additionally, we can reduce memory at inference time by constructing uncertainty directly in function space by linearizing the forward model (Foong et al., 2019; Immer et al., 2021). If the likelihood has the form $p(y\mid x,\theta)=p(y\mid f_{\theta}(x))$ and we have a posterior approximation $\operatorname{\mathbf{N}}(\theta\mid\mu,\Sigma)$ then the linearized forward distribution becomes $\operatorname{\mathbf{N}}(f(x)\mid f_{\mu}(x),\nabla_{\theta}f_{\mu}(x)^{\top}\Sigma\nabla_{\theta}f_{\mu}(x))$ where $\nabla_{\theta}f_{\mu}(x)$ is the Jacobian of $f_{\theta}(x)$ evaluated at $\mu$. This avoids having to sample many large parameter vectors.
@@
Stochastic gradient Markov chain Monte Carlo (SGMCMC, Ma et al. 2015; Nemeth and Fearnhead 2021) instead forms a Monte Carlo approximation $p(\theta\mid y_{1:N})\approx N^{-1}\sum_{k=1}^{K}\delta(\theta\mid\theta_{k})$. The samples $\{\theta_{k}\}_{k=1}^{K}$ are typically collected by (approximately) evolving a stochastic differential equation (SDE) - see Section 4 and Appendix B - and collecting samples along the trajectory. Traditional full-batch Markov chain Monte Carlo (MCMC) schemes (Andrieu et al., 2003) use a Metropolis-Hastings step to ensure convergence of the Monte Carlo approximation is preserved through discretization. In the minibatch setting, the Metropolis-Hastings ratio cannot be estimated easily. Instead, suppose we have access to an unbiased stochastic gradient. Then, as noted in Welling and Teh (2011), the bias in the trajectory decreases quadratically $\operatorname{\mathcal{O}}(\epsilon_{t}^{2})$ for learning rate $\epsilon_{t}$. Therefore, adopting a suitably decreasing learning rate schedule $\sum\epsilon_{t}^{2}<\sum\epsilon_{t}=\infty$ (Robbins and Monro, 1951) will ensure an asymptotically unbiased Monte Carlo approximation.
@@
  • Composability: posteriors is written in PyTorch (Paszke et al., 2019) and has compatibility with a broad range of tools including the Llama (Touvron et al., 2023) and Mistral (Jiang et al., 2023) models, lightning (Falcon and The PyTorch Lightning team, 2019) for convenient logging and device management, optimizers from torchopt (Ren et al., 2023) and probabilistic programs from pyro (Bingham et al., 2019), see Appendix G.
  • Extensible: The transform framework (Figure 3) adopted by posteriors is very general and allows for the easy adoption of new algorithms. Additionally, posteriors supports arbitrary likelihoods rather than being restricted to hard-coded regression or classification as is common in other libraries (Daxberger et al., 2021; Detommaso et al., 2023).
  • Functional: posteriors adopts a functional API via PyTorch’s (Paszke et al., 2019) functional module. As championed by the JAX (Bradbury et al., 2018) ecosystem, the functional approach makes for code that is easier to test and compose with other functions. Importantly for posteriors, functional programming is also closer to the mathematical description, which is particularly useful for Bayesian modelling.
  • Scalable: Unlike most common Bayesian software (Carpenter et al., 2017; Cabezas et al., 2024; Bingham et al., 2019), posteriors is minibatch-first, thus allowing for efficient computation in large datasets. Additionally, flexible sub-spacing is provided for scaling to large models.
@@
This combination of features makes posteriors unique amongst the PyTorch ecosystem which hosts the majority of the popular pretrained Hugging Face models (Wolf et al., 2020).
3.1 Supported Methods and Complexities
All methods discussed in Section 2 are included in posteriors. Flexible Laplace approximations are supported based on both the empirical Fisher and conditional Fisher (via the GGN matrix, see Section 2 and Appendix A). Laplace approximations rely on an optimization routine to find the MAP followed by a $O(bd)$ time and memory operation (for batch size $b$) to construct diagonal Fisher information matrices which ultimately require $O(d)$ memory. For very large models this can be a memory bottleneck meaning the batchsize may need to be reduced for the Laplace construction in comparison to the optimization (although the Laplace construction requires only a single epoch).
Variational inference for fitting a Gaussian distribution with diagonal covariance costs $O(kd)$ time and memory per iteration where $k$ is the number of samples from the variational distribution which is often set to 1 although larger $k$ reduces the variances of the gradient estimation. Dense covariance versions of both Laplace and VI are also supported but require $O(d^{2})$ memory and $O(d^{3})$ for sampling, however, this can still be useful for machine learning models when combined with subspace approaches (Mangrulkar et al., 2022; Daxberger et al., 2021).
As we will see in Section 4, SGMCMC is closely related to stochastic gradient descent and has the same per-iteration complexity. Although, a key difference is the final output is an ensemble of parameters rather than the single point estimate. As such memory requirements are $O(Md)$ where $M$ is the number of output samples, as with deep ensembles (Lakshminarayanan et al., 2017). Further, as a suitable decay in autocorrelation is desired between samples the number of iterations is typically longer for SGMCMC than optimization, although this can be parallelized, see Section 4.2.
In addition, posteriors hosts extended Kalman methods (Ollivier, 2018) (which has the same complexity as VI, Zhang et al. 2018), unified optimization via torchopt (Ren et al., 2023) and has a roadmap of methods to be added within the general framework including SVGD (Liu and Wang, 2016) and more advanced SGMCMC (Ma et al., 2015; Zhang et al., 2020). There is also a host of general-purpose utility functions not found elsewhere in the PyTorch ecosystem. This includes linearized forward calls, Fisher/GGN matrices, their diagonal counterparts, fast Fisher/GGN vector products and the conjugate gradient method, which provides the tools to implement Hessian-free optimization (Martens et al., 2010) and experiment with other second-order methods (Titsias, 2024; Duffield and Singh, 2022).
@@ 3.2 Limitations
posteriors supports arbitrary likelihoods and loss functions; as a result methods that utilise the specific form of the loss function (normally as classification or regression), e.g. the deterministic last layer approach in Harrison et al. (2024) or the probit approximations from Daxberger et al. (2021), are currently out of scope.
Similarly, the current methods in posteriors do not assume any specific structure of the model or its parameters, see Figure 3. However, this makes it more challenging to support model-aware techniques such as DVI (Wu et al., 2018) or K-FAC (Martens and Grosse, 2015), although the benefits of such approaches may warrant an inclusion in the library going forward. Specifically for K-FAC, the Fisher information matrix is assumed to exhibit a local block structure that admits cheap inversion and sampling via Kronecker-factors. Initially, K-FAC was only developed for feedforward layers (Martens and Grosse, 2015), before being extended to convolutional (Grosse and Martens, 2016) and recurrent (Martens et al., 2018) networks. More recently, it was generalized to weight-sharing networks (Eschenhagen et al., 2024) (including transformers and graph neural networks). Developing a flexible K-FAC submodule within posteriors’ functional framework is a high priority. Such a module should be suitably model-agnostic and capable of being applied across various inference methods (e.g., optimization, Laplace, VI). This advancement would enable scalable inference with non-diagonal Fisher information matrices, broadening the library’s capabilities.
Convergence diagnostics (e.g. Cowles and Carlin 1996) represent a key component of MCMC workflow, however are not included directly in posteriors due to the easy composition with Pyro’s (Bingham et al., 2019) extensive suite of convergence diagnostics. Similarly, posteriors is an inference library rather than a probabilistic programming language, therefore does not include tools such as distribution definitions which can be utilized through composition with Pyro as demonstrated in Appendix G. Finally, we opt not to include a fit convenience function (Chollet et al., 2015; Daxberger et al., 2021) in favour of the more flexible and transparent update framework in Figure 3.